OpenStudy (anonymous):

company a can install chairs in the theatre in 10 hours. company b can install them in 15 hours. the owner of a theatre wants the chairs installed in less than 8 hours. if the companies work together, can they install the chairs in less than 8 hours? justify.

OpenStudy (raden):

if the companies work together, the long time needed = 10*15/(10+15) = 150/25 = 6 yes, they are can

OpenStudy (anonymous):

thanks!=)

OpenStudy (raden):

welcome